Find the distinct interval of length 1 containing a root or solution of f(x)=x ³ - 3x + 5 using IVT
1
Expert's answer
2021-10-19T14:35:03-0400

Observe that

"f(-3)=(-3)^3-3(-3)+5=-13<0"

"f(-1)=(-1)^3-3(-1)+5=7>0"

Since "f" is continuous, we may conclude by the IVT that "f" has a root in "[-3, -1]."

Now


"f(-2)=(-2)^3-3(-2)+5=3>0"

So "f(-3)" and "f(-2)" are of opposite sign. Therefore, the IVT guarantees that "f" has a root on "[-3, -2]."

We find the interval of length 1 containing a root or solution of "f(x)=x^3 - 3x + 5."